# The Chicago-Carnegie Hubble Program:

### v2

@@ Abstract
The J-region asymptotic giant branch (JAGB) method is a new standard candle based on the constant luminosities of carbon-rich asymptotic giant branch stars in the J band. The JAGB method is independent of the Cepheid and TRGB distance indicators. Therefore, we can leverage it to both cross-check Cepheid and TRGB distances for systematic errors and use it to measure an independent local Hubble constant. The JAGB method also boasts a number of advantages in measuring distances relative to the TRGB and Cepheids, several of which are especially amplified when combined with JWST’s revolutionary resolving power. First, JAGB stars are 1 mag brighter in the NIR than the TRGB, and can be discovered from single-epoch NIR photometry unlike Cepheids which require congruent optical imaging in at least 12 epochs. Thus, JAGB stars can be used to measure significantly farther distances than both the TRGB stars and Cepheids using the same amount of observing time. Further advantages include: JAGB stars are easily identified solely via their colors and magnitudes, dust extinction is reduced in near-infrared observations, and JAGB stars are ubiquitous in all galaxies with intermediate-age populations. In this paper, we present a novel algorithm that identifies the optimal location in a galaxy for applying the JAGB method, so as to minimize effects from crowding. We then deploy this algorithm in JWST NIRCam imaging of seven SN Ia host galaxies to measure their JAGB distances, undertaking a completely blind analysis. The zero-point of this JAGB distance scale is set in the water mega-maser galaxy NGC 4258. In our CCHP $H_{0}$ overview paper Freedman et al. (2024), we apply the JAGB distances measured in this paper to the Carnegie Supernova Program (CSP) SNe Ia sample, measuring a Hubble constant of $H_{0}=67.96\pm 1.85~{}\rm{(stat)}\pm 1.90~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$.
@@ 1 Introduction
The CCHP compared TRGB and Cepheid distance moduli to 10 SN Ia host galaxies, finding a weighted averaged difference of $\rm{TRGB}-\rm{Cepheid}=+0.059$ mag, meaning the TRGB distances were measured to be systematically farther on average than the Cepheid distances. These significant differences indicate that investigating the potentially nefarious systematics of the local distance scale are still surely warranted before fully committing to the ‘new physics’ approach. Now, whether the Cepheid or TRGB distance ladder is more accurate in measuring distances is the primary question. One potentially powerful avenue to answer this is through utilizing a ‘tie-breaker’ distance ladder, i.e. comparing the TRGB and Cepheid SN Ia calibrator galaxy distances with distances measured from a third, independent yet equally accurate and precise distance indicator. However, until recently, such a pertinent distance indicator has yet to exist. That is, until the CCHP recently developed an incredibly promising new standard candle based on the constant luminosities of carbon-rich asymptotic giant branch (AGB) stars in the J band (1.2 microns), called the J-region asymptotic giant branch (JAGB) method.
In two seminal papers in 2020, we first proposed the JAGB method as a standard candle when we observed the J-band magnitudes of carbon stars were conveniently constant with zero color dependence, and consistent from galaxy to galaxy. We calibrated the JAGB method zeropoint using detached-eclipsing binaries in the LMC and SMC and then subsequently measured distances to 14 nearby galaxies (Madore & Freedman, 2020; Freedman & Madore, 2020), finding these distances agreed with TRGB distances at the 3% level. In the ensuing years, we have continued to test the JAGB method in preparation for a future $H_{0}$ measurement with JWST. For example, we have extensively shown the JAGB method is equally as precise and accurate at measuring distances as the TRGB and Cepheid Leavitt law in nearby galaxies using ground-based data (Lee et al., 2021, 2022, 2024b).
@@
The CCHP was fortunately awarded a JWST cycle 1 program (JWST GO 1995; PIs: W. L. Freedman, B. F. Madore), aimed at reducing the current systematics in the Cepheid, TRGB, and JAGB distance scales to provide our most accurate measurement of $H_{0}$ to date. This program, which measures JAGB, TRGB, and Cepheid distances to 10 SN Ia host galaxies, is described in the JWST CCHP $H_{0}$ overview paper (Freedman et al., 2024). The descriptions of our TRGB and Cepheid measurements of $H_{0}$ are also found in companion papers: Hoyt et al. (2024) and Owens et al. (2024), both in preparation.
@@ 2 Data
Observations were taken as part of the CCHP’s JWST cycle 1 program #1995 (PIs: W. Freedman, B. Madore), for which we imaged 11 galaxies (10 SN Ia hosts and 1 geometric anchor) from November 2022 to January 2024 with JWST’s NIRCam. NIRCam is JWST’s primary imager and can simultaneously observe in two channels, a short-wavelength (SW) channel and a long-wavelength (LW) channel, via a beam splitting dichroic which reflects the short wavelengths and transmits the long wavelengths. We designated the short-wavelength filter to be F115W (J band equivalent, $1.15$ microns) because the JAGB method must be applied as a standard candle in the J band, where the magnitudes of the JAGB stars are constant with color (Madore & Freedman, 2020). For the long-wavelength filter, we originally chose the reddest NIRCam long-wavelength filter F [^4] W ($4.42$ microns), to create the largest possible color baseline for separating our target carbon-rich AGB stars from contaminant populations like oxygen-rich AGB stars via their colors. However, after inspecting the images of the first two galaxies in our program, NGC 7250 and NGC 4536, we elected to switch our LW filter to F356W (3.56 microns) to take advantage of its increased angular resolution. We found the (F115W--F356W) color was as effective as the (F115W--F [^4] W) color at separating oxygen-rich and carbon-rich AGB stars. We kept F [^4] W as the LW filter in the two nearest galaxies, M101 and NGC 4258, to provide a metallicity test for our Cepheid program.222 The F [^4] W filter contains a CO bandhead that is sensitive to Cepheid metallicity (Scowcroft et al., 2016), so could theoretically be used to test for a radial Cepheid metallicity effect in M101 and NGC 4258.
@@ 2.1 Photometry
The details of the CCHP’s photometry procedure are described in Jang et al. (in prep). We briefly summarize them here. First, we acquired all level 2b *_cal.fits images from MAST. [^3] Our images were processed from the following JWST pipeline software versions: CAL_VER = 1.11.4, CRDS_VER = 11.17.2, SDP_VER = 2023_2a, and CRDS_CTX = jwst_1149.pmap. Next, we extracted PSF photometry from the images using the NIRCam module of DOLPHOT v2.0 (Dolphin, 2000, 2016; Weisz et al., 2023, 2024). We include the details of this procedure below. We also emphasize we performed this entire analysis blinded, by adding random offsets to the photometry which were only removed after the analysis was finalized.
First, we performed initial photometry on the individual images to detect the bright stars. We used these stars to re-align all the F115W images, and then drizzled them to create an astrometric reference image for DOLPHOT. Within DOLPHOT, we utilized the warmstart mode. Specifically, we extracted photometry solely from the SW F115W images, and then reduced the LW images using the source list positions of the stars from the first run. We found this procedure subtracted stars more cleanly from the images than running DOLPHOT on the SW and LW images simultaneously. We note the SHoES team (Riess et al., 2023) and a JWST TRGB calibration program (Newman et al., 2024) independently also found the warmstart mode resulted in the most cleanly subtracted residual images.
@@ 3.1 Description of JAGB Method’s Theoretical Basis
For AGB stars with masses of $\approx 2-5M_{\odot}$, a sufficient number of dredge-up events will eventually cause the abundance of carbon to exceed that of oxygen (C/O$>1$) on the stellar surface, and the star then transitions from an oxygen-rich to carbon-rich AGB star. This mass range occurs because AGB stars less massive than $\sim 2M_{\odot}$ fail to evolve into carbon stars since they lose their entire stellar envelope after a few thermal pulses (as the mass of the envelope was so small to begin with), and therefore transition into planetary nebula before the conversion to carbon star can take place. Stars more massive than $\sim 5M_{\odot}$ also unsuccessfully evolve into carbon stars because they undergo hot-bottom burning. Here, the carbon in the stars’ interior is burned into nitrogen as it is transported throughout the star because the star is so massive and therefore extremely hot (Habing & Olofsson, 2003; Herwig, 2005; Marigo et al., 2013). Therefore, only AGB stars with a narrow range of initial masses (and therefore luminosities) eventually evolve into carbon stars. Thus, the small range of NIR magnitudes observed for JAGB stars ($M_{J}=-6.2\pm 0.3$ mag) can be straightforwardly attributed to the astrophysics of AGB stars. Herein lies the foundational theoretical basis for carbon stars as standard candles.
Carbon stars can also be easily photometrically distinguished from other stellar populations because the $\rm{C_{2}}$ and CN molecules in their atmospheres increase their opacity in typical photometric bandpasses. Therefore, carbon stars have cooler effective temperatures (and therefore much redder colors) relative to their oxygen-rich predecessors (Marigo et al., 2003). We can thus separate carbon stars from oxygen-rich AGB stars solely via their near-infrared colors. In Figure 3, we show a NIR CMD of one of the SN Ia host galaxies from our program, NGC 4639. The carbon stars can be cleanly delineated from other stellar populations via their colors alone. Furthermore, contamination from competing stellar populations is almost negligible, because JAGB stars are the brightest and reddest stellar populations in a given galaxy ($M_{J}=-6.2$ mag). Only background galaxies lie in the same color-magnitude space as JAGB stars, but these can be straightforwardly eliminated using sharpness cuts (see Section 2.1).
@@ 3.3 Advantages and Disadvantages of the JAGB Method
A possible limitation of the JAGB method is that theoretically, metallicity is predicted to influence the JAGB stars’ magnitudes. Theory predicts the minimum mass for carbon star formation is larger at higher metallicity, because the third dredge-up event’s efficiency decreases with increasing metallicity (Pastorelli et al., 2020). Furthermore, carbon stars in metal-rich environments contain more oxygen in their atmospheres, which will preferentially bind with the dredged-up carbon to form CO instead of carbon molecules such as CN or $\rm{C_{2}}$, thereby also hindering the carbon-enhancement of these stars’ atmospheres. Thus, metallicity should theoretically influence the shape and mode of the carbon star LF.
@@ 3.4 Measuring JAGB distances to SN Ia Host Galaxies
Here, we summarize the CCHP’s procedure for measuring JAGB distances. First, we color selected the JAGB stars. For galaxies with F [^4] W data, JAGB stars were selected as having colors between $3.22.4<(F115W-F444W)<$ 3.2 mag. For galaxies with F356W data, JAGB stars were selected as having colors between $3.02.3<(F115W-F356W)<$ 3.0 mag. Then, we binned the F115W magnitudes of the JAGB stars using bins of 0.01 mag. Next, we smoothed the binned luminosity function using a nonparametric interpolation technique: the Gaussian-windowed, Locally Weighted Scatterplot Smoothing (GLOESS) algorithm. The GLOESS smoothing technique is effective at suppressing false (noise-induced) edges, and has been used in several astrophysical contexts like smoothing variable-star light curves (e.g., Persson et al. 2004) and for smoothing the RGB LF for measuring the TRGB (e.g., Freedman et al. 2019). The only user input is the smoothing parameter $\sigma_{s}$. For all galaxies, we used a smoothing parameter of $\sigma_{s}=0.25$ mag. In Section 3.6.1, we describe how we incorporated an uncertainty due to our choice of smoothing parameter. The peak location (mode) of the smoothed luminosity function then marks the JAGB magnitude.
@@ 3.4.1 Choice of Spatial Selection for the JAGB method
    Next, we numerically determined $dm/dr$, the derivative of $m_{JAGB}$ as a function of radial distance.
@@
We also noticed the presence of a spiral arm in the convergence plots sometimes caused additional noisiness. This caused $m_{JAGB}$ to converge at a radial distance with too few JAGB stars or $dm/dr$ to never equal zero in four galaxies: M101, NGC 2442, NGC 4258, and NGC 4639. We masked these spiral arms, and then re-calculated the convergence plots, where $m_{JAGB}$ then successfully converged. The convergence plots without the spiral arms masked are shown in grey in Figure 4. We emphasize we performed this procedure during the blinded stage of our analysis. We only masked the spiral arms so that $m_{JAGB}$ successfully converged, not to change the final measured distance. Leaving the spiral arms unmasked while using the newly adopted radial cuts yielded almost a negligible change in $H_{0}$ of 0.3% (larger). This is also further discussed in Appendix C.
We observed the same pattern in all our convergence plots where the mode was generally measured to be brightest in the inner, high-surface brightness regions of a galaxy. In Appendix A, we show CMDs for the ‘inner region’ of each galaxy, which were produced using the stars inside the chosen radial cut. In our final sample of seven calibrating galaxies, we found the average difference between the measured JAGB magnitude in the inner vs. outer regions to be $\rm{<JAGB_{inner}-JAGB_{outer}}>=-0.21$ mag, meaning the JAGB magnitude was on average 0.21 mag brighter in the inner disks of these galaxies. Thus, we continue to caution the JAGB magnitude can be significantly biased when measured in the inner disks of galaxies. This pattern may result from crowding, the effect of which will be larger in the higher surface brightness regions in galaxies.

### v3

@@ Abstract
The J-region asymptotic giant branch (JAGB) method is a new standard candle based on the constant luminosities of carbon-rich AGB stars in the J band. The JAGB method is independent of the Cepheid and TRGB distance indicators. Therefore, we can leverage it to both cross-check Cepheid and TRGB distances for systematic errors and use it to measure an independent local $H_{0}$. The JAGB method also boasts a number of advantages in measuring distances relative to the TRGB and Cepheids, several of which are especially amplified when combined with JWST’s revolutionary resolving power. First, JAGB stars are 1 mag brighter in the NIR than the TRGB, and can be discovered from single-epoch NIR photometry unlike Cepheids which require congruent optical imaging in at least 12 epochs. Thus, JAGB stars can be used to measure significantly farther distances than both the TRGB stars and Cepheids using the same amount of observing time. Dust extinction is also reduced in near-infrared observations and JAGB stars are ubiquitous in all galaxies with intermediate-age populations. We present a novel algorithm that identifies the optimal location in a galaxy for applying the JAGB method, so as to minimize effects from crowding. We then deploy this algorithm in JWST NIRCam imaging of seven SN Ia host galaxies to measure their JAGB distances, undertaking a completely blind analysis. In our CCHP $H_{0}$ results paper Freedman et al. (2025), we apply the JAGB distances measured in this paper to the Carnegie Supernova Program (CSP) SNe Ia sample, measuring a Hubble constant of $H_{0}=67.80\pm 2.17~{}\rm{(stat)}\pm 1.64~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$.
@@ 1 Introduction
In Freedman et al. (2019), the CCHP compared TRGB and Cepheid distance moduli to 10 SN Ia host galaxies, finding a weighted averaged difference of $\rm{TRGB}-\rm{Cepheid}=+0.059$ mag, meaning the TRGB distances were measured to be systematically farther on average than the Cepheid distances. These significant differences indicate that investigating the potentially nefarious systematics of the local distance scale are still surely warranted before fully committing to the ‘new physics’ approach. Now, whether the Cepheid or TRGB distance ladder is more accurate in measuring distances is the primary question. One potentially powerful avenue to answer this is through utilizing a ‘tie-breaker’ distance ladder, i.e. comparing the TRGB and Cepheid SN Ia calibrator galaxy distances with distances measured from a third, independent yet equally accurate and precise distance indicator. However, until recently, such a pertinent distance indicator has yet to exist. That is, until the CCHP recently developed an incredibly promising new standard candle based on the constant luminosities of carbon-rich asymptotic giant branch (AGB) stars in the J band (1.2 microns), called the J-region asymptotic giant branch (JAGB) method.
In two seminal papers in 2020, Madore & Freedman (2020); Freedman & Madore (2020) first proposed the JAGB method as a standard candle when they observed the J-band magnitudes of carbon stars were conveniently constant with zero color dependence, and consistent from galaxy to galaxy. Madore & Freedman (2020); Freedman & Madore (2020) then calibrated the JAGB method zeropoint using detached-eclipsing binaries in the LMC and SMC and then subsequently measured distances to 14 nearby galaxies, finding these distances agreed with TRGB distances at the 3% level. In the ensuing years, the CCHP has continued to test the JAGB method in preparation for a future $H_{0}$ measurement with JWST. For example, we have extensively shown the JAGB method is equally as precise and accurate at measuring distances as the TRGB and Cepheid Leavitt law in nearby galaxies using ground-based data (Lee et al., 2021, 2022, 2024b). The JAGB method has also been extensively tested by several other groups (e.g., Ripoche et al. 2020; Zgirski et al. 2021; Parada et al. 2021, 2023; Li et al. 2024).
@@
The CCHP was fortunately awarded a JWST cycle 1 program (JWST GO 1995; PIs: W. L. Freedman, B. F. Madore), aimed at reducing the current systematics in the Cepheid, TRGB, and JAGB distance scales to provide our most accurate measurement of $H_{0}$ to date. This program, which measures JAGB, TRGB, and Cepheid distances to 10 SN Ia host galaxies, is described in the JWST CCHP $H_{0}$ results paper (Freedman et al., 2025). The descriptions of our TRGB and Cepheid measurements of $H_{0}$ are also found in companion papers: Hoyt et al. (2025) and Owens et al. (in preparation).
@@ 2 Data
Observations were taken as part of the CCHP’s JWST cycle 1 program #1995 (PIs: W. Freedman, B. Madore), for which we imaged 11 galaxies (10 SN Ia hosts and 1 geometric anchor) from November 2022 to January 2024 with JWST’s NIRCam. A more detailed description of our observing program can be found in Hoyt et al. (2024). NIRCam is JWST’s primary imager and can simultaneously observe in two channels, a short-wavelength (SW) channel and a long-wavelength (LW) channel, via a beam splitting dichroic which reflects the short wavelengths and transmits the long wavelengths. We designated the short-wavelength filter to be F115W (J band equivalent, $1.15$ microns) because the JAGB method must be applied as a standard candle in the J band, where the magnitudes of the JAGB stars are constant with color (Madore & Freedman, 2020). For the long-wavelength filter, we originally chose the reddest NIRCam long-wavelength filter F [^4] W ($4.42$ microns), to create the largest possible color baseline for separating our target carbon-rich AGB stars from contaminant populations like oxygen-rich AGB stars via their colors. However, after inspecting the images of the first two galaxies in our program, NGC 7250 and NGC 4536, we elected to switch our LW filter to F356W (3.56 microns) to take advantage of its increased angular resolution. We found the (F115W--F356W) color was as effective as the (F115W--F [^4] W) color at separating oxygen-rich and carbon-rich AGB stars. We kept F [^4] W as the LW filter in the two nearest galaxies, M101 and NGC 4258, to provide a metallicity test for our Cepheid program.222 The F [^4] W filter contains a CO bandhead that is sensitive to Cepheid metallicity (Scowcroft et al., 2016), so could theoretically be used to test for a radial Cepheid metallicity effect in M101 and NGC 4258.
@@ 2.1 Photometry
The details of the CCHP’s photometry procedure are described in Jang et al. (in prep). We briefly summarize them here. First, we acquired all level 2b *_cal.fits images from MAST. [^3] Our images were processed from the following JWST pipeline software versions: CAL_VER = 1.11.4, CRDS_VER = 11.17.2, SDP_VER = 2023_2a, and CRDS_CTX = jwst_1149.pmap. Next, we extracted PSF photometry from the images using the NIRCam module of DOLPHOT v2.0 (Dolphin, 2000, 2016; Weisz et al., 2023, 2024). We include the details of this procedure below. We also emphasize we performed this entire analysis blinded. That is, we performed the entire analysis without knowing the final measured distances. In practice, we added random photometric offsets to the photometry which were only removed after the analysis was finalized in Section 3.4.1. Our blinding procedure is described in detail in the CCHP $H_{0}$ results paper in Freedman et al. (2025).
First, we performed initial photometry on the individual images to detect the bright stars. We used these stars to re-align all the F115W images, and then drizzled them to create an astrometric reference image for DOLPHOT. Within DOLPHOT, we utilized the warmstart mode. Specifically, we extracted photometry solely from the SW F115W images, and then reduced the LW images using the source list positions of the stars from the first run. We found this procedure subtracted stars more cleanly from the images than running DOLPHOT on the SW and LW images simultaneously. We note the SHoES team (Riess et al., 2023) and a JWST TRGB calibration program (Newman et al., 2024) independently also found the warmstart mode resulted in the most cleanly subtracted residual images. All our photometry is on the Sirius-Vega calibration system.
@@ 3.1 Description of JAGB Method’s Theoretical Basis
For AGB stars with masses of $\approx 2-5M_{\odot}$, a sufficient number of dredge-up events will eventually cause the abundance of carbon to exceed that of oxygen (C/O$>1$) on the stellar surface, and the star then transitions from an oxygen-rich to carbon-rich AGB star. This mass range occurs because AGB stars less massive than $\sim 2M_{\odot}$ fail to evolve into carbon stars since they lose their entire stellar envelope after a few thermal pulses (as the mass of the envelope was so small to begin with), and therefore transition into planetary nebula before the conversion to carbon star can take place. Stars more massive than $\sim 5M_{\odot}$ are also thwarted from evolving into carbon stars because they undergo hot-bottom burning. Here, the carbon in the stars’ interior is burned into nitrogen as it is transported throughout the star because the star is so massive and therefore extremely hot (Habing & Olofsson, 2003; Herwig, 2005; Marigo et al., 2013). Therefore, only AGB stars with a narrow range of initial masses (and therefore luminosities) eventually evolve into carbon stars. Thus, the small range of NIR magnitudes observed for JAGB stars ($M_{JAGB}~{}\pm~{}0.3$ mag) can be straightforwardly attributed to the astrophysics of AGB stars. Herein lies the foundational theoretical basis for carbon stars as standard candles.
Carbon stars can also be easily photometrically distinguished from other stellar populations because the $\rm{C_{2}}$ and CN molecules in their atmospheres increase their opacity in typical photometric bandpasses. Therefore, carbon stars have cooler effective temperatures (and therefore much redder colors) relative to their oxygen-rich predecessors (Marigo et al., 2003). We can thus separate carbon stars from oxygen-rich AGB stars solely via their near-infrared colors. In Figure 3, we show a NIR CMD of one of the SN Ia host galaxies from our program, NGC 4639. The carbon stars can be cleanly delineated from other stellar populations via their colors alone. Furthermore, contamination from competing stellar populations is almost negligible, because JAGB stars are the brightest and reddest stellar populations in a given galaxy ($M_{JAGB}=-6.2$ mag in the ground-based J band). Only background galaxies lie in the same color-magnitude space as JAGB stars, but these can be straightforwardly eliminated using sharpness cuts (see Section 2.1).
@@ 3.3 Advantages and Disadvantages of the JAGB Method
A possible limitation of the JAGB method is that theoretically, metallicity is predicted to influence the JAGB stars’ magnitudes. Theory predicts the minimum mass for carbon star formation is larger at higher metallicity, because the third dredge-up event’s efficiency decreases with increasing metallicity (Pastorelli et al., 2020). Furthermore, carbon stars in metal-rich environments contain more oxygen in their atmospheres, which will preferentially bind with the dredged-up carbon to form CO instead of carbon molecules such as CN or $\rm{C_{2}}$, thereby also hindering the carbon-enhancement of these stars’ atmospheres. Thus, the mode of the JAGB LF metallicity is theorized to be brighter in metal-rich environments. (Pastorelli et al., 2020)
@@ 3.4 Measuring JAGB distances to SN Ia Host Galaxies
Here, we summarize the CCHP’s procedure for measuring JAGB distances. First, we color selected the JAGB stars. For galaxies with F [^4] W data, JAGB stars were selected as having colors between $3.22.4<(F115W-F444W)<$ 3.2 mag. For galaxies with F356W data, JAGB stars were selected as having colors between $3.02.3<(F115W-F356W)<$ 3.0 mag. Then, we binned the F115W magnitudes of the JAGB stars using bins of 0.01 mag. Next, we smoothed the binned luminosity function using a nonparametric interpolation technique: the Gaussian-windowed, Locally Weighted Scatterplot Smoothing (GLOESS) algorithm (Cleveland & Loader, 1996; Loader, 2004; Persson et al., 2004). The GLOESS smoothing technique is effective at suppressing false (noise-induced) edges, and has been used in several astrophysical contexts like smoothing variable-star light curves (e.g., Persson et al. 2004) and for smoothing the RGB LF for measuring the TRGB (e.g., Freedman et al. 2019). The only user input is the smoothing parameter $\sigma_{s}$. For all galaxies, we used a smoothing parameter of $\sigma_{s}=0.25$ mag. In Section 3.6.1, we describe how we incorporated an uncertainty due to our choice of smoothing parameter. The peak location (mode) of the smoothed luminosity function then marks the JAGB magnitude.
@@ 3.4.1 Choice of Spatial Selection for the JAGB method
    Next, we numerically determined $dm/dr$, the smoothed derivative of $m_{JAGB}$ as a function of radial distance. [^9] The code used to calculate $dm/dr$ can be found at https://github.com/abiglee7/CCHP-JAGB/blob/main/jagb_der.py.We smoothed $dm/dr$ according to the same criteria for each galaxy, where the smoothing interval was 1/6 the number of $m_{JAGB}$ data points.
@@
    Furthermore, to prevent selecting a radial cut where the derivative equaled zero due to random fluctuations, we imposed an additional criteria. We calculated the maximum $m_{JAGB}$, minimum $m_{JAGB}$, median $m_{JAGB}$, and standard deviation $\sigma$ of the $m_{JAGB}$ data points outside of the selected radial cut. The maximum and minimum $m_{JAGB}$ were then required to be greater and less than the median±$\pm 2\sigma$, respectively. The smoothing interval was increased until this criterion was fulfilled. This criterion affected three galaxies, M101, NGC 2442, and NGC 4258, where the first zero-crossing of the smoothed derivative was clearly due to a random fluctuation and not where the derivative actually converged.
@@
We also noticed the presence of a spiral arm in the convergence plots sometimes caused additional noisiness. This caused $m_{JAGB}$ to converge at a radial distance with too few JAGB stars or $dm/dr$ to never equal zero in four galaxies: M101, NGC 2442, NGC 4258, and NGC 4639. We masked these spiral arms using increasingly larger masks until $m_{JAGB}$ successfully converged. The functional forms of the masking criteria are detailed in Appendix A. The convergence plots without the spiral arms masked are shown in grey in Figure 4. We emphasize we performed this procedure during the blinded stage of our analysis. We only masked the spiral arms so that $m_{JAGB}$ successfully converged, not to change the final measured distance. Leaving the spiral arms unmasked while using the newly adopted radial cuts yielded almost a negligible change in $H_{0}$ of 0.3% (larger). This is also further discussed in Appendix A.
We observed the same pattern in all our convergence plots where the mode was generally measured to be brightest in the inner, high-surface brightness regions of a galaxy. In Appendix B, we show CMDs for the ‘inner region’ of each galaxy, which were produced using the stars inside the chosen radial cut. In our final sample of seven calibrating galaxies, we found the average difference between the measured JAGB magnitude in the inner vs. outer regions to be $\rm{<JAGB_{inner}-JAGB_{outer}}>=-0.21$ mag, meaning the JAGB magnitude was on average 0.21 mag brighter in the inner disks of these galaxies. Thus, we continue to caution the JAGB magnitude can be significantly biased when measured in the inner disks of galaxies. This pattern may result from crowding, the effect of which will be larger in the higher surface brightness regions in galaxies.